OpenStudy (anonymous):

What is log(x − 2)(2x − 3) = logx 2

OpenStudy (solomonzelman):

`(x - 2)` is not the base, right ?

OpenStudy (anonymous):

Right

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \log (x − 2)(2x − 3) = \log x^2 }\) like this ?

OpenStudy (anonymous):

Yah

OpenStudy (solomonzelman):

Okay, so we will expand the first log, can you expand `(x-2)(2x-3)` ?

OpenStudy (anonymous):

To (x-2)(2x-3) = 2x^2-7x+6 ?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Mmkay

OpenStudy (solomonzelman):

And now we have \(\normalsize\color{blue}{ \log (2x^2-7x+6) = \log (x^2) }\). We know however, that when we have \(\normalsize\color{red}{ \log( a )= \log (b) }\) , then we are able to conclude that \(\normalsize\color{red}{ a = b }\). So in our case, `2x² - 7x+6` is "a", and `x²` is "b" .

OpenStudy (solomonzelman):

\(\normalsize\color{darkgoldenrod}{ \log( 2x^2-7x+6)= \log (x^2) }\) turns into ?

OpenStudy (anonymous):

There Are Two Solutions x = 6 and x = 1 Right ?